# Applications and resource estimates for open system simulation on a quantum computer

### v1

@@ Abstract
We study applications of a fully controllable open system quantum simulator. A universal quantum computer can realize such a simulator, and some classical methods can also approximate it. We introduce two concrete computational problems, such that finding their solution would have direct scientific or industrial utility. The scientific utility is exemplified by a computation of nonequilibrium behavior of Ca3Co2O6, which has been studied in the costly MagLab experiments. Specifically, an order of \$2M per material can be saved if an affordable quantum computer simulation is used to screen the materials before sending them to MagLab. For industrial utility, we develop a methodology that allows researchers of various backgrounds to estimate the economic value of an emerging technology consistently. We then apply our approach to the applications of materials with a Metal-Insulator Transition. These materials have not been used commercially yet, but many potential applications, including alternative transistors, neuromorphic computing, and smart windows, amount in total to \$20M for the entire material search performed on a quantum computer. We provide zeroth-order resource estimates for both algorithms on superconducting qubit hardware, finding that simulating long-lifetime nonequilibrium effects presents a new challenge for quantum simulators. Finally, we introduce planted solution problems and their obfuscated versions to test the capabilities of the future quantum device. The intended use of those problems is to match the size of application benchmarks so that solving one guarantees the ability to solve the other.
@@ I Introduction
We study applications of an open system quantum simulator. Such a simulator can be realized by a universal quantum computer [cleve2016efficient], and some classical methods can approximate it as well. Our goal is to introduce a concrete computational problem such that finding its solution would have direct scientific or industrial utility. We find an example of the scientific utility to be the investigation of materials that are already studied in high-cost experiments. 20-30 new material samples per year are not understood by low-cost experimental and computational means and are sent to the Magnetic Field Laboratory (MagLab) [national2013high]. These experiments often observe unexplained nonequilibrium phenomena, exemplified by the behavior of Ca3Co2O6 [nekrashevich2022reaching]. We use them as application instances.
@@
For industrial utility, we focus on material searches. A successful example of a large throughput material search is the search for material for good thermoelectric devices, happening under the umbrella of the Materials Project[jain2013commentary]. There are already commercial applications for those devices, but wide adoption is still dependent on the discovery of better materials and designs using the semi-automated search methodology. The utility of this computational search comes from accelerating their adoption and the projected big markets for applications. Unfortunately for quantum computers, we believe the bottlenecks in this search are largely classical, and most questions can and should be answered using cheaper state-of-the-art classical algorithms. Drawing inspiration from the success of these classical techniques, we envision a similar search in the domain of strongly correlated materials. There, classical methods fail, though there are some promising new results [cui2023ab]. The commonly mentioned applications for these materials include high-temperature superconductivity and metal-insulator transition. The latter is especially suited for an open system study, as one of the central observables is the transport coefficient. It can, in principle, be estimated in a closed system setting, yet we believe the most efficient quantum algorithm for transport contains an open system quantum simulator. The corresponding economic impact will be outlined below.
@@
We also note that many other application instances covered in companion publications, while not directly related to open systems, have their solution by quantum algorithms relying on an unproven assumption of the existence of a high overlap initial state. Estimating that overlap is a challenge, and here we note that for open system thermal state preparation relevant to most condensed matter and material science applications, as well as in quantum chemistry such as surface chemistry, the estimate can be performed under very mild assumptions compared to other state preparation methods.
@@
In Sec. II, we discuss our general metodology of utility estimation and apply it to cases of interest. Then in Sec III we state two of the cases of interest as fully specified mathematical problems. We list available quantum algorithms in Sec. IV, before describing implementation of the best of them for our problems in Sec. V, calculating zeroth order estimates to both logical and physical required resources. Finally, in App. A we use toy problem benchmarks where we can compute the answer to test the capabilities of the future quantum device. The intended use of those planted solution benchmarks is to match the size of the application benchmarks so that solving one guarantees the ability to solve the other.
@@ II Utility
There are two approaches to the utility of scientific findings: to look into the funding science receives and to look into the market of inventions enabled by the discovery in question. The latter is only possible in hindsight, while the former may be insensitive to differences between fundamental and applied research. In an efficient market, one would expect both estimates to result in a similar order of magnitude.
We note that the quantum computer is not necessarily the only way to accomplish the research agenda we outline. Whether it would be accomplished by classical computation, experiment, quantum computer, or a combination of those approaches, the utility is unchanged. In particular, the presence or absence of quantum advantage should not affect our estimate of utility. Nonetheless, we focus on the problems that promise a quantum advantage as well. Below, we will outline the methodology of each approach and attempt both approaches for the chosen application instances.
@@ II.1 Scientific utility: funding
For the funding estimate, we look into the amount of money spent on solving the problem in question or a problem of equivalent size and significance. The rationale for that is that one capable of solving the problem should be able to negotiate for that amount of remuneration, at least in principle. We note that a transaction like that would not necessarily occur. For example, the creators of AlphaFold did not monetize it in a way commensurate with the amount of money regularly spent on protein structure experiments and simulations. Possibly, the publicity it generated was of equivalent value to them, and it undoubtedly accelerated biological research, so we believe estimating its value is still appropriate using our methodology even if the hypothetical transaction did not take place.
@@ II.2 General methodology for industrial utility
The argument that the particular technology’s market size is the utility of further research into it is very common. Almost any research is conventionally justified in that way, no matter how distant it is from applications. We would like to challenge this way of thinking by proposing a methodology that would result in a consistent (within 1-2 orders of magnitude) estimate of utility even when applied by different researchers and decision makers independently. While it is true that it is sometimes impossible to trace the influences of fundamental research on applications far in the future, most of the revolutionary technologies of the past have had clear conditions for their discovery. We choose to focus on concrete links between research and applications. Thus, only the research that is part of a clearly defined roadmap to an improvement in the corresponding technology will be assigned nonzero utility in this approach. In this way, we will estimate the effort required for the successful completion of that roadmap and assign the utility to that.
@@
We will now describe how to utilize each of those sources.
@@
Our approach may fail in more complex situations such as when one of the technologies has already received mass adoption. In that case, while there is a probability that the entire industry will switch to a new technology, complete switches like that are rare. E.g., no one is building coal plants in the US anymore, but the existing ones are still in use, and sometimes their operation is even extended beyond the planned closure date. The switch may not be complete and will have a lot of inertia. The available data does not allow us to reliably estimate this switch probability - indeed, it is hard to formulate a methodology that will result in the same (within 1-2 orders of magnitude) estimate when applied independently by different analysts or research groups. Instead, we rely on a more tangible picture that the new technology will occupy a niche of the market where it is especially advantageous.
@@
The data for such market segmentation is usually available in the industry reports published by consulting companies. Those reports can be rather controversial, such as the trillion-dollar estimates of the utility of quantum computers themselves. Below, we outline the methodology we use to interpret those reports, as well as the precautions that need to be taken when using them.
The market size of a technology is commonly estimated as the sum of the yearly revenue of the companies utilizing that technology. More specifically, if a small part of a big company is using it, the corresponding share is used. The revenue is the product of the product’s price and the number of sales. We note that technology is definitely not enough to create a product; one also needs people and assets, as well as various other costs. While a company cannot fill the market in question without the corresponding technology, it would not be accurate to count the entire revenue as the utility of the technology. The founder who discovers the technology either invests his own money or seeks venture funding by selling shares of the company when the company is created. Only with that money can the company fill the market. Since there are various monetary figures at various stages of this process, and the revenue itself grows with time, we need a standardized method to reduce it all to a single number - utility. Note that the first stage of venture funding usually values the company relatively cheap, no matter how good the underlying technology is. We see that as the tradition of venture capital and not as the true value of the corresponding technology. We believe that the true value of technologies shows itself in acquisition events, where a big company acquires promising ideas before they become competitors. We also note that for utility estimates, the total revenue is less important than the relative increase with or without the technology. Ideally, that can be computed from the business model and an estimate of how much the technology improves the product. That being said, none of those numbers can easily be used in a unified framework, as they are rarely available. For simplicity, we will assign a factor between $10$ (technology will lead to an explosive growth of the business) and $1/10$ depending on how central the technology is for the business model. For example, if energy is the leading cost and improvement in efficiency directly translates to energy savings, such as for nitrogen fixation, we choose the factor $1/2$. A more detailed estimate of that factor would be as follows: first, we assume a simple model of seller optimizing profit, and the number of sales n depending linearly on price as $n=n_{0}-ep$, where e is the elasticity of demand. The result of profit optimization for production costs $c1$ and $c2$ is that the difference in revenue is $r_{2}-r_{1}=\frac{1}{2}e(c_{1}-c_{2})$. As the fertilizer market is tied with the population and the arable land, we can estimate elasticity as the remaining need for fertilizer divided by the current price since, at zero price, only the amount needed for current land and population would be purchased. The amount of fertilizer needed would remain approximately the same over a 10-year span. Finally, we estimate lower production costs $c_{2}$ from the energy consumption to fix nitrogen. Various methods have $O(1)$ difference in energy cost per unit of mass of nitrogen fixed, so we can expect a novel method to have of order 50% improvement. If it is only a 1% improvement, the estimate can be updated by directly multiplying the factors. Assuming the 50% energy savings, prices close to the production cost, and elasticity $\sim n_{0}/$current price, this estimate results in a factor $1/2$ improvement in revenue for the new nitrogen fixation technology.
In contrast, a slightly better product requiring a lot of investment into switching the manufacturing process completely would have two competing effects: rising production costs would lower the revenue, yet the improvement in product quality would increase the revenue, though the effect may be quite complex to formulate. For example, consider thermoelectric devices. While they are sold commercially today, the market is small, mostly consisting of niche cooling applications. The main application of recycling heat is mostly unrealized. It is believed that improvements in efficiency (essentially, the percentage of heat recycled) would make it viable and result in mass adoption. More specifically, the efficiency, together with other figures of merits involving lifespan, production, and maintenance cost of the thermoelectric device, determines the viability for various use cases. The demand in each use case would discontinuously jump from zero to mass adoption, but if one adds these jumps for all use cases, one could approximate the number of sales as linear in the figures of merit. Using the formalism above, that would be $n_{0}=aZT$. The change in revenue would then be $en_{0}a(ZT_{2}-ZT_{1})$. If one expects a tenfold increase in the number of sales, that would roughly translate to a tenfold increase in revenue, assuming other coefficients are $O(1)$. That illustrates that technologies with a small existing market can have much higher utility than the current or projected market size suggests. This applies only to conventionally impossible tasks, e.g. the profitable heat recycling is essentially impossible with the current technology, therefore the existing market is not a good estimate for the potential market. For most technologies, however, the task is already accomplished by the older generation of technology, and the shift in demand is not as dramatic, no matter how good the new generation of technology is.
@@
A related question is, at what time should we take the revenue? With the quantum computer application in mind, we need to wait for at least ten years before the solution to the problem becomes available. Most projections of the market size stop at the 10-year time horizon. Thus, we will use an estimate of the market size ten years in the future (year 2034) corresponding to the emerging technology or competing technology in case the market is already occupied by existing technology. Unfortunately, the market breakdown based on emerging technologies is not commonly found in consulting reports. At best, we can narrow it down to the niche of the existing technology to be replaced. What is common is another type of estimate: instead of looking at the revenue of the seller of the technology, a report would look at the revenue of all the potential buyers of the technology. Then, the report would note that an improvement in the product would generate some arbitrary fraction, such as a 1%-9% increase in the revenue of the buyers. The possibility of such an improvement is confirmed by experts in the buyers’ corresponding businesses. Thus, the total value is computed as a sum of 1%-9% of the revenues of all buyers, and the resulting number for some of the future technologies, such as AI or quantum computing, is quoted in the trillions. An immediate objection is that the revenue of the current semiconductor industry is just \$500b global and \$80b in the US. If they really had a trillion dollar value, they would have been able to increase their revenue correspondingly. Indeed, the consulting message for the semiconductor industry is that they only capture 20% of the value their product generates, and it is the goal for the CEOs of the hardware companies to increase that ratio for future hardware markets, such as specialized AI hardware. For our purposes, we will only look at the value the seller companies were able to realize as revenue, not the optimistic possibility that is quoted in the more controversial reports. Moreover, only when a company based in one country can, in principle, take over the entire global market with a superior technology that cannot be reverse-engineered we use the global revenues instead of the revenue of companies based in one particular country.
@@
A final caution is that a ”buzzword” technology attracts scam artists. As a result, only scam industry reports can be found for the current market revenue of some ”buzzword” technologies. For example, looking for the market size of neuromorphic devices, one only finds scam reports. A well-known consulting company’s reports are available for the dedicated AI hardware but not for the neuromorphic devices. They likely have not generated enough revenue to report. If we aim to replace dedicated AI hardware with neuromorphic devices, we will need to use the dedicated AI hardware market size and estimate the probability of the switch in some way. The estimate of the market size of a segment like that may vary by two orders of magnitude between reports. Whenever possible, we average the reports of well-known consulting companies, thus achieving one order of magnitude error. As the other factors we used are ad hoc, this estimate would result in two orders of magnitude error. The formula is as follows:

### v2

@@ Abstract
We present two applications where open system quantum simulation is the preferred approach on a quantum computer. We choose concrete parameters for the problems in such a way that the application value, which we call utility, can be obtained from the solution directly. The scientific utility is exemplified by a computation of nonequilibrium behavior of Ca3Co2O6, which is studied in \$2M MagLab experiments. For industrial utility, we develop a methodology that allows researchers of various backgrounds to estimate the economic value of an emerging technology consistently. Our approach predicts \$400M utility for the applications of materials with a Metal-Insulator Transition. We focus on the transport calculation in the Hubbard model as the simplest problem that needs to be solved in a large-scale material search. The resource estimates for both problems suffer from a large required runtime, which motivates us to propose novel algorithm optimizations, taking advantage of the translation invariance and the parallelism of the T-gate application. Finally, we introduce several planted solution problems and their obfuscated versions as a benchmark for future quantum devices.
@@ I Introduction
We study applications of an open system quantum simulator. Such a simulator can be realized by a universal quantum computer [cleve2016efficient], and some classical methods can approximate it as well. Our goal is to introduce a concrete computational problem such that finding its solution would have direct scientific or industrial utility. We find an example of the scientific utility to be the investigation of materials that are already studied in high-cost experiments. 20-30 new material samples per year are not understood by low-cost experimental and computational means and are sent to the National High Magnetic Field Laboratory (MagLab) [national2013high]. These experiments often observe unexplained nonequilibrium phenomena, exemplified by the behavior of Ca3Co2O6 [nekrashevich2022reaching]. We use them as application instances.
@@
For industrial utility, we focus on material searches. A successful example of a large throughput material search is the search for material for good thermoelectric devices, happening under the umbrella of the Materials Project [jain2013commentary]. There are already commercial applications for those devices, but wide adoption is still dependent on the discovery of better materials and designs using the semi-automated search methodology. The utility of this computational search comes from accelerating their adoption and the projected big markets for applications. Unfortunately for quantum computers, we believe the bottlenecks in this search are largely classical, and most questions can and should be answered using cheaper state-of-the-art classical algorithms. Drawing inspiration from the success of these classical techniques, we envision a similar search in the domain of strongly correlated materials. There, classical methods fail, though there are some promising new results [cui2023ab]. The phenomena that are most often discussed in the context of applications are high-temperature superconductivity and the metal-insulator transition (MIT). The latter is especially suited for an open system study, as one of the central observables is the transport coefficient. It can, in principle, be estimated in a closed system setting, yet we believe the most efficient quantum algorithm for transport contains an open system quantum simulator. The corresponding economic impact will be outlined below. The majority of the MIT search will need to investigate ab initio models similar to those considered in [cui2023ab]; however, as a starting point, we choose a simpler Hubbard model as a lower bound of the required computational resources of this approach.
@@
We also note that many other application instances covered in companion publications [watts2024fullereneencapsulatedcyclicozonegeneration, otten2024quantumresourcesrequiredbinding, nguyen2024quantumcomputingcorrosionresistantmaterials, bellonzi2024feasibilityacceleratinghomogeneouscatalyst, penuel2024feasibilityacceleratingincompressiblecomputational, elenewski2024prospectsnmrspectralprediction, agrawal2024quantifyingfaulttolerantsimulation, bartschi2024potentialapplicationsquantumcomputing, saadatmand2024faulttolerantresourceestimationusing], while not directly related to open systems, are expected to be solved by quantum algorithms that rely on an unproven assumption of the existence of a high overlap initial state. Estimating that overlap is a challenge, and here we note that for open system thermal state preparation relevant to most condensed matter and material science applications, as well as in quantum chemistry such as surface chemistry, the estimate can be performed under very mild assumptions compared to other state preparation methods.
In this work, we perform resource estimates for quantum algorithms solving our chosen open system simulation problems for Ca3Co2O6 and the Hubbard model. The complexity of the best open system quantum simulation algorithms is only known in terms of big-O notation, which limits our estimates as we ignore the prefactors. Prefactors can be readily estimated for suboptimal quantum simulation. While simulating the high-temperature Hubbard model can be performed within a realistic time scale of hours, low-temperature ab initio simulations appear unfeasible without further algorithmic improvements. An even simpler complication we encounter is that for any material like Ca3Co2O6, where nonequilibrium dynamics takes weeks of real time, the ratio between the microscopic time scale and the timescale of the slow dynamics is $10^{11}$ or higher. This ratio directly multiplies the required algorithm runtime, resulting in unrealistic numbers of the order of a billion years. The only ways around this issue are fast-forwarding of the dynamics or utilizing NISQ quantum simulators such as D-wave [king2021quantum]. As the runtime appears to be the deciding factor in the practicality of these algorithms, we adopt an unusual hardware compilation scheme. The conventional approach [beverland2022assessing, harrigan2024expressing] gets rid of all the Clifford operations, which forces it to perform remaining T-gates sequentially. Our estimates and algorithm optimizations utilizing translation invariance take advantage of the parallel application of T-gates. This can be realized in hardware if Clifford gates are performed explicitly. The CNOTs that are not spatially local, in particular, need to be performed via logical qubit teleportation, where entangled Bell pairs become a resource consumed throughout the computer in a similar way to magic states. This approach has a high physical qubit count, which is the tradeoff for better runtime.
@@
In Sec. II we discuss our general methodology of utility estimation and apply it to cases of interest. The relevant economic concepts are further discussed in App. A. Then, in Sec III, we state two of the cases of interest as fully specified mathematical problems. We list available quantum algorithms in Sec. IV before describing the implementation of the best of them for our problems in Sec. V, calculating zeroth order estimates for both logical and physical required resources. In App. B we use toy problem benchmarks where we can compute the answer to test the capabilities of the future quantum device. The intended use of those planted solution benchmarks is to match the size of the application benchmarks so that solving one guarantees the ability to solve the other. App. C describes how to count T-gates and depth for the operations used in our algorithm and App. D provides estimates for a concrete computational problem that is a proxy for the average case of the MIT material search (in contrast to Hubbard, which was the easiest case). App. E describes how we utilized the pyLIQTR [pyLIQTR] library for the block encoding of the Hubbard model Hamiltonian, and in App. F we compute the exact resource estimates for a simpler suboptimal algorithm to contrast it with the order of magnitude estimates used in the main text.
@@ II Utility
There are two approaches to the utility of scientific findings: to look into the funding that science receives and to look into the market of inventions enabled by the discovery in question. The latter is possible only in hindsight, while the former may be insensitive to differences between fundamental and applied research. For private research, one may assume an efficient market, which would require both estimates to result in a similar order of magnitude. The efficiency of government-funded research depends on the policy of the corresponding funding agency. We expect it to be at least within an order of magnitude of efficiency.
We note that the quantum computer is not necessarily the only way to accomplish the research agenda we outline. Whether it would be accomplished by classical computation, experiment, quantum computer, or a combination of those approaches, the utility is unchanged. In particular, the presence or absence of quantum advantage should not affect our estimate of utility. Nonetheless, we focus on the problems that promise a quantum advantage as well. In the following, we will outline the methodology of each approach to utility estimation and try both approaches for the chosen application instances.
@@ II.1 Scientific utility: funding
For the funding estimate, we look into the amount of money spent on solving the problem in question or a problem of equivalent size and significance. The rationale for that is that one capable of solving the problem should be able to negotiate for that amount of remuneration, at least in principle. We note that a transaction like that would not necessarily occur. For example, the creators of AlphaFold [Alphafold2021] did not monetize it in a way commensurate with the amount of money regularly spent on protein structure experiments and simulations. Possibly, the publicity it generated was of equivalent value to them, and it undoubtedly accelerated biological research, so we believe estimating its value is still appropriate using our methodology even if the hypothetical transaction did not take place.
There is a subtle difference between estimating the utility of an answer to one fully specified computational problem, and the utility of a black box solving any problem in some general class. We focus on the former, while [agrawal2024quantifyingfaulttolerantsimulation] establish a data-driven methodology for the latter, using the publication data in the corresponding field.
@@ II.2 General methodology for industrial utility
The argument that a particular technology’s market size is the utility of further research into it is very common. Almost any research is conventionally justified in that way, no matter how distant it is from applications. We would like to challenge this way of thinking by proposing a methodology that would result in a consistent (within 1-2 orders of magnitude) estimate of utility even when applied by different researchers and decision-makers independently. While it is true that it is sometimes impossible to trace the influences of fundamental research on applications far in the future, most of the revolutionary technologies of the past have had clear conditions for their discovery. We choose to focus on concrete links between research and applications. Thus, only the research that is part of a clearly defined roadmap to an improvement in the corresponding technology will be assigned nonzero utility in this approach. In this way, we will estimate the effort required for the successful completion of that roadmap and assign the utility to that.
@@
We will now describe how to utilize these sources.
@@
Our approach may fail in more complex situations such as when one of the technologies has already received mass adoption. In that case, while there is a probability that the entire industry will switch to a new technology, complete switches like that are rare. E.g., no one is building coal plants in the US anymore, but the existing ones are still in use, and sometimes their operation is even extended beyond the planned closure date [coal]. The switch may not be complete and will have a lot of inertia. The available data does not allow us to reliably estimate this switch probability - indeed, it is hard to formulate a methodology that will result in the same (within 1-2 orders of magnitude) estimate when applied independently by different analysts or research groups. Instead, we rely on a more tangible picture that the new technology will occupy a niche of the market where it is especially advantageous.
@@
The data for such market segmentation is usually available in the industry reports published by consulting companies. Those reports can be rather controversial, such as the trillion-dollar estimates of the utility of quantum computers themselves [bogobowicz2024steady]. Below, we outline the methodology we use to interpret those reports, as well as the precautions that need to be taken when using them.
The market size of a technology is commonly estimated as the sum of the yearly revenue of the companies utilizing that technology. More specifically, if a small part of a big company is using it, the corresponding share is used. The revenue is the product of the product’s price and the number of sales. For simplicity, we will assign a factor between $10$ (technology will lead to an explosive growth of the business) and $1/10$ depending on how central the technology is for the business model. We discuss the possible considerations going into this number further in App. A. For example, if energy is the leading cost and improvement in efficiency directly translates to energy savings, such as for nitrogen fixation, it is reasonable to choose the factor $1/2$. For thermoelectric devices, the factor is $10$: an improvement in efficiency and other qualities may dramatically increase the market, unlocking the main application of energy recycling.
@@
A related question is, at what time should we take the revenue? With the quantum computer application in mind, we need to wait for at least ten years before the solution to the problem becomes available. Most projections of the market size stop at the 10-year time horizon. Thus, we will use an estimate of the market size ten years in the future (year 2034) corresponding to the emerging technology or competing technology in case the market is already occupied by existing technology. Unfortunately, the market breakdown based on emerging technologies is not commonly found in consulting reports. At best, we can narrow it down to the niche of the existing technology to be replaced. What is common is another type of estimate: instead of looking at the revenue of the seller of the technology, a report would look at the revenue of all the potential buyers of the technology. Then, the report would note that an improvement in the product would generate some arbitrary fraction, such as a 1%-9% increase in the revenue of the buyers. The possibility of such an improvement is confirmed by experts in the buyers’ corresponding businesses. Thus, the total value is computed as a sum of 1%-9% of the revenues of all buyers, and the resulting number for some of the future technologies, such as AI or quantum computing, is quoted in the trillions [bogobowicz2024steady]. The apparent discrepancy with the \$600b total semiconductor market is intentional: the message of those reports is that the semiconductor companies could generate more revenue, pushing the software middlemen out and realizing their full potential. This message is likely biased: it is a sales pitch for the semiconductor companies to hire the consulting services of the company that made the report. For our purposes, we will only look at the value the seller companies were able to realize as revenue, not the optimistic possibility that is quoted in the more controversial reports. Moreover, only when a company based in one country can, in principle, take over the entire global market with a superior technology that cannot be reverse-engineered can we use global revenues instead of the revenue of companies based in one particular country.
@@
A final caution is that fake consulting reports exist. Especially for emerging technologies such as neuromorphic computing, all the reports available online are of dubious origin and quote seemingly random numbers as the market size. A well-known consulting company’s reports are available for the dedicated AI hardware but not for the neuromorphic devices. The companies invested in them likely have not generated enough revenue to report. Assuming that the dedicated AI hardware of the future will use neuromorphic devices, we can use the dedicated AI hardware market size and estimate the probability of the switch in the way outlined above. The estimate of the market size of a segment like that may vary by two orders of magnitude between reports. Whenever possible, we average the reports of well-known consulting companies, thus achieving one order of magnitude error. As the other factors we used are ad hoc, this estimate would result in two orders of magnitude error. The formula is as follows: